OpenStudy (4everaddicted2anime):

Find the difference in the volume and total area of a cylinder with both a radius and height of 1. r = 1, h = 1 The number of sq units of the total area exceeds the number of cu. units in the volume by 3pi 2pi pi

rebeccaxhawaii (rebeccaxhawaii):

well do you have the formula?

OpenStudy (4everaddicted2anime):

No

OpenStudy (4everaddicted2anime):

What is the formula?

OpenStudy (fortytherapper):

\[V = \pi*r^2*h\] \[A = 2*\pi*r^2*+ 2*\pi*r*h\]

OpenStudy (fortytherapper):

They said sq exceeds cm, so \[Area-Volume\]

OpenStudy (pawanyadav):

Total area of cylinder= curve surface area + area of circular end = 2Πrh + 2Πr^2

OpenStudy (4everaddicted2anime):

Okay so for volume I would have V= 1^2 *1 * pi V=pi And for area I would have A=2pi 1^2 + 2pi 1 * 1 A=2pi + 2pi A=4pi 4pi - pi= 3pi Is this right?

OpenStudy (fortytherapper):

Looks good to me

OpenStudy (4everaddicted2anime):

Okay. Thank you